OpenStudy (anonymous):

A bag contains 1 blue, 2 white, and 2 red marbles. Match the columns. Drag the letter after a compound event in the left column and drop it before the matching probability in the right column. Term: P(red and red), with replacement P(white and white), with replacement P(blue and red), with replacement P(white and red), with replacement Definition: A) 20% B) 16% C) 8% D) 10%

OpenStudy (anonymous):

mach each term with a Definition

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

you need these probabilities yes?

OpenStudy (anonymous):

5 marbles in the bag. red and red with replacement \[\frac{2}{5}\times \frac{2}{5}=\frac{4}{25}=.16=16\%\]

OpenStudy (anonymous):

one min plz

OpenStudy (anonymous):

back

OpenStudy (anonymous):

white and white same as red and red

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

blue and red \[2\times \frac{1}{5}\times \frac{2}{5}=16\%\]